OpenStudy (christina166):

I will fan and medal !!! Prove that the two circles shown below are similar.

OpenStudy (christina166):

@pooja195 Please, please help I have been stuck with this for hours.

OpenStudy (bobo-i-bo):

It easily follows from the mathematical definition of being "similar". What is that defintion?

OpenStudy (christina166):

Of similar?

OpenStudy (bobo-i-bo):

What does it mean that two circles are similar?

OpenStudy (christina166):

It means that they are both equal to the fixed point

OpenStudy (bobo-i-bo):

No that is that is not the definition of two objects being similar: http://www.mathsisfun.com/geometry/similar.html

OpenStudy (christina166):

Oh okay, so what do I put to proof that the circles are similar?

OpenStudy (bobo-i-bo):

Two objects are similar if you can resize, rotate, reflect and translate rotate one of the objects into the other

OpenStudy (christina166):

Yep, thank you, but how do I prove it using those two circles?

OpenStudy (bobo-i-bo):

So you have to prove that, after a combination of resizing, rotation, reflection and translation, you can place one circle directly on top of the other. Does that make sense?

OpenStudy (christina166):

Yes, but I don't know how to prove it. D:

OpenStudy (christina166):

What I mean is you explained it but with the coordinates of the circle I don't know how to prove.

OpenStudy (bobo-i-bo):

So, let's try and place the little circle directly over the bigger circle. In order to do that, i need the little circle the same size and centered exactly where the bigger circle is. That would prove that they're similar

OpenStudy (christina166):

Oh okay, so how do I make them the same size and centered exactly?

OpenStudy (bobo-i-bo):

So it basically comes down to the question of: How many units do I need to shift the centre of the little circle upwards and leftwards to get it to be at the centre of the bigger circle? And then how much do i need to scale the little circle by to make it the same size as the bigger circle?

OpenStudy (bobo-i-bo):

sorry, that should say rightwards and not leftwards

OpenStudy (christina166):

You need to dilate it.

OpenStudy (bobo-i-bo):

To know how much you need to scale the little circle, consider the radius of the little circle and the radius of the bigger circle

OpenStudy (christina166):

Oh okay.

OpenStudy (christina166):

So its (3,4) (0,-1)

OpenStudy (bobo-i-bo):

ummm, yes, you need to move the centre from (0,-1) to (3,4) so how many units do you need to shift it from (0,-1) to (3,4)?

OpenStudy (bobo-i-bo):

*how many units rightwards and upwards do you need to shift it from (0,-1) to (3,4)?

OpenStudy (christina166):

(3,5)

OpenStudy (bobo-i-bo):

:)

OpenStudy (bobo-i-bo):

So in proper terminology, we say that we are applying a translation of 3 units rightwards and 5 units upwards

OpenStudy (christina166):

Okay. Then what do we do (By the way, thanks so much I'm actually understanding)

OpenStudy (bobo-i-bo):

If you would like to get a bit more technical, then we are applying a translation vector \[\left(\begin{matrix}3 \\ 5\end{matrix}\right)\]. Tells you about it here: http://www.bbc.co.uk/schools/gcsebitesize/maths/geometry/transformationsrev2.shtml

OpenStudy (bobo-i-bo):

Okay, so now we have translated it and have the little circle's centre right where the bigger circle's centre is, we need to scale the little circle up so that we can get the little circle directly over the bigger circle

OpenStudy (christina166):

Okay

OpenStudy (bobo-i-bo):

Following me? :P Any idea on how to do it?

OpenStudy (christina166):

Yep but I just don't know how to do it next

OpenStudy (bobo-i-bo):

okay, so what is the radius of the little circle and what is the radius of the bigger circle?

OpenStudy (christina166):

Um by let me see.

OpenStudy (christina166):

By 2pir

OpenStudy (bobo-i-bo):

sorry, i don't understand :P what is 2pir?

OpenStudy (christina166):

Sorry I ment x=r2/r1

OpenStudy (bobo-i-bo):

Yup! :D

OpenStudy (bobo-i-bo):

Actually, i've just realised a slight complication :P

OpenStudy (christina166):

So how do I plug it in D:

OpenStudy (bobo-i-bo):

oh... let's deal with that first

OpenStudy (bobo-i-bo):

so do you know what is r2 and r1?

OpenStudy (christina166):

Radius 1 and Radius 2?

OpenStudy (bobo-i-bo):

And what are those values?

OpenStudy (christina166):

I don't know :/

OpenStudy (bobo-i-bo):

Radius 1 is the radius of the little circle, and radius 2 is the radius of the bigger circle. So what are the radii of the little and bigger circles?

OpenStudy (christina166):

I don't know. ://

OpenStudy (christina166):

Can you tell me but show me how you got it so I learned

OpenStudy (bobo-i-bo):

Radius 1 is the radius of the little circle, and radius 2 is the radius of the bigger circle. So what are the radii of the little and bigger circles?

OpenStudy (christina166):

How do I get the redii

OpenStudy (bobo-i-bo):

The length of the red line is the length of radius 1 and length of green line is the length of radius 2

OpenStudy (christina166):

How do I calculate it though?

OpenStudy (bobo-i-bo):

So, the endpoints of the green line are (-2, 4) and (3,4) so the length of the green line is?

OpenStudy (bobo-i-bo):

And the length of the green line IS radius 2

OpenStudy (christina166):

(5, 0)

OpenStudy (bobo-i-bo):

Well, the length of the green line is 5

OpenStudy (christina166):

THE LENGTH IS 2 FOR THE RED LINE

OpenStudy (bobo-i-bo):

:D

OpenStudy (christina166):

So then what ://

OpenStudy (bobo-i-bo):

So r1=2, r2=5 and then in order to scale the little circle to the big circle, we need ti scale it by a scale factor of r2/r1

OpenStudy (christina166):

it equals to 2.5

OpenStudy (bobo-i-bo):

Indeed :)

OpenStudy (christina166):

So whats after that

OpenStudy (bobo-i-bo):

Lol, so what have u done so far?

OpenStudy (christina166):

I have

OpenStudy (christina166):

First you have to apply a translation of 3 units rightwards and 5 units upwards. Which is a translation vector. Then you use x=r2/r1 which the bigger circle equals (5,0) and the smaller equals (2,0) this equals to 2.5.

OpenStudy (bobo-i-bo):

Hmmm, so the translation thing is nice and clear

OpenStudy (bobo-i-bo):

The rest isn't as clear :P

OpenStudy (christina166):

How do I fix it, may you help :(

OpenStudy (bobo-i-bo):

For the second part we wish to say this: We apply a dilation of scale factor ___ about the centre (__,__) to the smaller circle

OpenStudy (bobo-i-bo):

You need to fill in the blanks :P

OpenStudy (bobo-i-bo):

Does that make sense?

OpenStudy (christina166):

Okay and the blank space for the first one is x=r2/r1 and the second blank space is 2.5?

OpenStudy (bobo-i-bo):

No, x=r2/r2=5/2=2.5

OpenStudy (bobo-i-bo):

x IS the scale factor

OpenStudy (bobo-i-bo):

the scale factor IS 2.5 which you calculated by dividing r2 by r1

OpenStudy (christina166):

for the first one I put r2/r1 and for the second I put 5/2=2.5

OpenStudy (bobo-i-bo):

nono, they are both the same thing haha

OpenStudy (christina166):

Oh yesss so for the first I put 5/2= 2.5

OpenStudy (bobo-i-bo):

yes :)

OpenStudy (christina166):

And for the second I put (2,5)

OpenStudy (bobo-i-bo):

no, not (2,5)

OpenStudy (bobo-i-bo):

it should be (3,4) because that is where the centre of little circle is.

OpenStudy (christina166):

Ohh okay, so I have First you have to apply a translation of 3 units rightwards and 5 units upwards. Which is a translation vector. I then apply a dilation of the scale factor 5/2=2.5 about the center (3,4) to the smaller circle.

OpenStudy (bobo-i-bo):

You don't need to mention about the translation vector thing. Don't worry about it.

OpenStudy (bobo-i-bo):

Apart from that, yes! :D

OpenStudy (bobo-i-bo):

That's the proof

OpenStudy (christina166):

YAYAYAYA. so then I put at the end "Which proves that the two circles will be similar"?

OpenStudy (bobo-i-bo):

Yes, because as we have proved, after some combination of dilation, rotation, translation and reflection, we can make one circle completely identical to the other. ^_^

OpenStudy (christina166):

Oh okay thank you so much!!!!

OpenStudy (bobo-i-bo):

np ^_^

OpenStudy (owen3):

.